Berta Company recently lost its entire inventory in a fire. The following information is available from its accounting records: Beginning inventory: $1,000; purchases: $13,000; net sales: $20,000. The company's average gross profit percentage is 40%. Using the gross profit method, a reasonable estimate of the lost inventory would be
Answer #1
  • Inventory cost at beginning : $1000
  • Purchase : $13,000
  • Sales : $20000
  • cost of Goods Available = $1000 + $13,000 = $14,000
  • Gross Profit percentage is 40%. So Cost of Goods Sold = 100-40 = 60%
  • Cost of Goods Sold = $20000 * 60% = $12000
  • Ending Inventory = Cost of Goods Available - Cost of Goods Sold = $14000 - $12000 = $2000

Lost Inventory would be $2000